OpenStudy (anonymous):

Suppose y varies directly with x, and y = 25 when x = 140. What is the value of x when y = 36? (1 point)] 140 165 201.6 176

OpenStudy (anonymous):

Ratio of x/y = 140/25. Now, multiply whatever value of y by this ratio and you will get value of x corresponding to the given value of y. In this case, you get 36 * 140/25 = 201.6

OpenStudy (anonymous):

Suppose y varies directly with x, and y = 14 when x = –4. What is the value of y when x = –6? (1 point) –21 10 18 21 what wpuld this one be? @GT

OpenStudy (anonymous):

For your second question the answer is 21